OpenStudy (anonymous):

the 2 stands for "squared" in the following equation: g(t) = at2-a2t find g(-1/2) and g(a)

OpenStudy (bahrom7893):

So a(t^2) - (a^2)t ?

OpenStudy (bahrom7893):

Helloo? If that's right, then g(-1/2)=a(-1/2)^2-(a^2)(-1/2) = (a/4)+(a^2/2)

OpenStudy (anonymous):

wait im trying to figure out how to review the whole chat lol

OpenStudy (anonymous):

i guess i have to hit post every time?

OpenStudy (bahrom7893):

yea. These are like threads.

OpenStudy (anonymous):

ohh okay, so basically i just plug everything in? sorry im very new to this, thanks for your patience

OpenStudy (bahrom7893):

yea g(a) -> replace all ts with as g(-1/2) -> replace all ts with (-1/2)

OpenStudy (anonymous):

how did you get 0 g(a)

OpenStudy (bahrom7893):

g(a) = a*a^2-a^2*a=a^3-a^3 = 0

OpenStudy (bahrom7893):

Now you'll have to hover your mouse over my profile pic and click on Become a fan, and add a testimonial describing whether I was helpful or not. And also click on the Best Response button if my response helped.

OpenStudy (swissgirl):

hahahahah

OpenStudy (anonymous):

ill do anything you want for helping me lol. i dont have nayone to help me out with this stuff so if you dont mind if like to keep comin back

OpenStudy (anonymous):

next time ill be better prepared. right now we're learning ranges of domain and functions of F

OpenStudy (bahrom7893):

lol ofc not. And if I'm not around, there're plenty of others who can help you out

OpenStudy (bahrom7893):

by the way if you need to ask a new question, you will need to close this one, on the left where you asked, there's a close button.

OpenStudy (bahrom7893):

And then you'll be able to ask a new question. If your question is lost below all the new ones, you can hit bump, and that will bump your question up to the top.

OpenStudy (bahrom7893):

Good luck and enjoy OS

OpenStudy (anonymous):

thanks but hon that top problem, ow did you get ato the third power?

OpenStudy (anonymous):

a^3 *

OpenStudy (bahrom7893):

a^2*a = a^2*a^1=a^(2+1)=a^3

OpenStudy (bahrom7893):

same base, different powers. When you multiply two numbers, you add their powers (if they have the same base)

OpenStudy (anonymous):

hah! i understand

OpenStudy (anonymous):

so basically boils down to a^3-a^3, right?

OpenStudy (bahrom7893):

yea